OpenStudy (anonymous):

Solve for x 3x - 6 = 7x + 18

OpenStudy (hockeychick23):

3x-3x-6=7x-3x+18 -6= 4x+18 -6-18= 4x -24= 4x -24/4= 4x/4 -6=x

OpenStudy (anonymous):

thnx